# PET: Preference Evolution Tracking with LLM-Generated Explainable Distribution

### v2

@@ Abstract
Understanding how user preference evolves over time is a fundamental challenge central to modern digital ecosystems, for which Large Language Models (LLMs) are an increasingly prominent and popular approach due to their ability to comprehend the rich semantic context within behavioral data. A common practice is to use LLMs to predict a user’s next action by directly generating a ranked list of preferred items. Although effective for short-term prediction, the end-to-end generation paradigm inherently limits personalization. Its opaque decision-making process obscures holistic user profiling and exacerbates popularity bias. To address these limitations, we propose Preference Evolution Tracking (PET), a framework that reframes the task as inferring a dynamic probability distribution over a stable and interpretable lattice of preference clusters. By applying logit-probing and generative classification techniques, PET infers a user’s preference as a probability distribution, enabling transparent preference learning. On public benchmarks (Yelp, MovieLens), PET improves ranking quality by up to $40\%$ in NDCG over direct generation baselines. On a large-scale, real-world dataset from a short-video platform, it excels at ranking long-tail contents, significantly outperforming a SOTA production model by $7$ times in the NDCG score. Ultimately, PET transforms the user profile model from direct preference list generation to a transparent distributional preference mapping, paving the way for more explainable, fair, and diverse personalization systems.
@@ 1 Introduction
Understanding user interests in a transparent and interpretable manner is a central challenge in building trusted and adaptive digital ecosystems. The advent of Large Language Models (LLMs) presents a transformative opportunity in this domain. With their profound ability to comprehend context and semantics within unstructured user behavioral histories – from product reviews to content consumption sequences – LLMs promise to move beyond simple pattern matching to a deeper, more holistic understanding of user intent. The ultimate goal is not merely to predict the next action of users, but to create transparent and dynamic models of user interest that can be trusted, analyzed, and fairly acted upon.
The prevailing approach for capturing user preferences involves leveraging LLMs to directly generate ranked outputs (wang2025user; wang2024llms; ngo2024recgpt; yu2024icpl; deng2025onerec), a method proven effective for short-horizon, top-$k$ recommendation tasks. While effective for short-horizon prediction and identifying popular preferences, this direct-to-ranking paradigm leaves important gaps. First, it lacks interpretability: although LLMs encode rich internal states, these models do not expose an explicit or structured representation of the user’s overall preferences – making them difficult to audit or reuse in downstream tasks. Second, the focus on a small $k$, combined with popularity bias, often leads to the over-representation of mainstream content – neglecting the long-tail interests that define a user’s niche tastes and are critical for achieving holistic, unbiased personalization.
These limitations have significant consequences for algorithmic decision-making in user-oriented applications. The lack of an explicit user model undermines system transparency, making it difficult to govern and audit for fairness and bias. Additionally, the over-reliance on popular items not only distorts model predictions but also reinforces popularity during training – suppressing the model’s ability to learn users’ nuanced, long-tail preferences. This feedback loop amplifies the Matthew Effect (i.e., the rich get richer), reinforcing echo chambers that reduce content diversity, entrench mainstream tastes, and marginalize underrepresented creators and user segments (bakshy2015exposure; wu2024survey). This motivates a paradigm shift toward transparent, distributional user modeling, leading to our central research question:
How can we infer personalized evolving preference distribution via LLM – enhancing interpretability, fairness, and diversity?
To address this problem, we propose Preference Evolution Tracking (PET), a framework that customizes LLMs into probabilistic inference engines for modeling users’ evolving preference distributions. By leveraging pre-softmax logits over a stable preference cluster lattice (e.g., categories or tags), PET captures structured relevance signals more effectively than volatile item-level spaces. This approach enhances interpretability and diversity, especially for long-tail preferences. Specifically, PET supports two scalable inference strategies: Likelihood-based Probing, which iteratively queries the model for each cluster to build a precise distribution, and the more efficient Generative Classification, which extracts all probabilities in a single forward pass. Furthermore, PET employs Hierarchical Probing to maintain tractable inference in extreme multi-label scenarios through semantic taxonomy traversal. Theoretically, the ranking according to the inferred PET distribution guarantees an optimal ranking results (e.g., with respective to NDCG/Recall) under standard assumptions. Applied to domain-aligned LLMs, PET produces portable, interpretable user profiles for long-horizon ranking, fairness auditing, and long-tail preference modeling.
To validate PET’s effectiveness, we conduct comprehensive experiments across three diverse datasets with varying cluster complexity: MovieLens, Yelp, and a large-scale dataset from a world-leading short video platform. Our results demonstrate that the distributional approach significantly outperforms direct-generation baselines and state-of-the-art (SOTA) ranking models on public benchmarks. More critically, on the real-world data of the short video platform, our framework shows a unique ability to rank users’ niche, long-tail interests, significantly outperforming a SOTA model in production. Notably, our goal is not to replace mature production models on highly popular (head) content. Rather, PET complements them by recovering user-specific long-tail preferences that are typically underserved by popularity-optimized systems. Across all experiments, our methods prove more adept at capturing stable, long-term preferences over transient, short-term interests. Our primary contributions are threefold:
@@
    A New Paradigm for User Modeling: We introduce a novel framework that shifts the focus from optimizing short-term item predictions to generating interpretable, dynamic probability distributions over a stable lattice of human-understandable preference clusters.
@@
    A Portable and Dynamic User Profile: The learned preference distribution serves as a rich, multi-faceted user profile that captures long-tail interests and is readily portable for diverse downstream tasks, such as user segmentation, trend analysis, or fairness audits.
@@
    Theoretical and Empirical Validation: We prove that PET’s distributional ranking is theoretically superior to direct generation under standard assumptions. Empirically, we validate PET on three datasets (MovieLens, Yelp, short-video platform), achieving $40\%$ NDCG over direct generation and $20\%$ NDCG over a SOTA ranking model (Qwen3-Reranker-8B (qwen3embedding)) on public benchmarks, and a $7\times$ improvement in long-tail ranking against a SOTA production-level ranking baseline on a real-world dataset.
Figure 1: The PET framework pipeline. Left (Training): An LLM is trained on user history to learn preference distributions. Center (Inference): probing methods extracts the predicted preference distribution from the model’s internal logits. Right (Application): The transparent distribution is used for downstream tasks: personalized ranking, long-tail discovery, and interpretable user profiling.
@@ 2 Related Work
User Preference Modeling. To deliver effective personalization, user models aims to capture preferences that are both diverse and dynamic, through three key lines of work. Multi-interest models represent users via multiple embeddings linked to distinct interests (shi2023everyone; cen2020controllable; zhou2018deep). Sequential models, such as FPMC (rendle2010factorizing), GRU4Rec (hidasi2015session), SASRec (kang2018self), and BERT4Rec (sun2019bert4rec), focus on temporal dynamics to predict the next item in a user’s history. More recently, distributional modeling has become central in alignment and fairness, with methods like GDPO for group fairness (yao2024no) and RLHF-based techniques like DPL and AOT for reward uncertainty (siththaranjan2023distributional; melnyk2024distributional). While powerful, these alignment-based methods introduce significant practical considerations, as they often require costly preference-labeled datasets and complex, computationally intensive pipelines.
LLMs in Preference Modeling. Recent work has increasingly applied Large Language Models (LLMs) to the complex task of user preference modeling. Early efforts leveraged LLMs as powerful components for embedding generation (U-BERT; qiu2021u) or unified the field under a prompt-based, text-to-text paradigm (P5; geng2022recommendation). This led to end-to-end generative models like OneRec (deng2025onerec) and (ngo2024recgpt) that directly produce ranked outputs or the most relevant items (wang2025user; wang2024llms), often incorporating sophisticated alignment modules. In a parallel effort toward explainability, RecGPT uses the LLM to generate descriptive, textual user profiles (yi2025recgpt). Crucially, the direct generation of ranked lists via LLM – the approach that forms the basis of our primary baseline – has been shown to consistently outperform traditional sequential models (like SASRec (kang2018self) and BERT4Rec (sun2019bert4rec)) in various settings (geng2022recommendation; wu2024survey; zhao2024recommender; deng2025onerec).
Recent approaches also tackle complex scenarios through structured prompting and multi-context modeling, such as LLM4MSR (wang2024llm4msr), HUM (bao2025heterogeneous), and user interest exploration via latent clusters (wang2024llms). While effective for short-horizon prediction, these methods typically encode a user’s state implicitly within the model’s parameters, without surfacing an interpretable or temporally grounded preference distribution. Our work is also related to logit-probing and generative classification techniques for extracting model beliefs from pre-softmax logits (petroni2019language; schick2020s; raffel2020exploring; zhou2024quest), adapting them to infer evolving user preference distributions from fine-tuned LLMs.

### v3

# SPECTRA: Revealing the Full Spectrum of User Preferences via Distributional LLM Inference

@@ Abstract
Large Language Models (LLMs) are increasingly used to understand user preferences, typically via the direct generation of ranked item lists. However, this end-to-end generative paradigm inherits the bias and opacity of autoregressive decoding, over-emphasizing frequent (head) preferences and obscure long-tail ones, thereby biasing personalization toward head preferences. To address this, we propose SPECTRA (Semantic Preference Extraction and Clustered TRAcking), which treats the LLM as an implicit probabilistic model by probing it to infer a probability distribution over interpretable preference clusters. In doing so, SPECTRA reframes user modeling from sequence generation with decoding heuristics to distributional inference, yielding explicit, cluster-level user preference representations. We evaluate SPECTRA on MovieLens, Yelp, and a large-scale short-video platform, demonstrating significant gains across three dimensions: SPECTRA achieves (i) distributional alignment, reducing Jensen–Shannon divergence to empirical distributions by $25\%$ against strong baselines; (ii) long-tail exposure, reducing decoding-induced head concentration and increasing global exposure entropy by $30\%$; and (iii) downstream applications such as personalized ranking, translating these gains into a $40\%$ NDCG boost on public datasets and a 7× improvement on ranking long-tail preferences against an industry-leading Transformer-based production baseline.
SPECTRA: Revealing the Full Spectrum of User Preferences via Distributional LLM Inference
Luyang Zhang thanks: Part of this work was done while at TikTok Inc. Carnegie Mellon University luyangz@andrew.cmu.edu          Jialu Wang TikTok Inc. faldict@ucsc.edu          Shichao Zhu TikTok Inc. shichao.szhu@gmail.com
Beibei Li Carnegie Mellon University beibeili@andrew.cmu.edu          Zhongcun Wang TikTok Inc.          Guangmou Pan TikTok Inc.
Yang Song TikTok Inc. ys@sonyis.me
@@ 1 Introduction
Understanding user preferences in an explicit and interpretable manner is a central challenge in building personalized AI systems (Doshi-Velez and Kim, 2017; Amershi et al., 2019; Rudin, 2019; Lipton, 2018). Large Language Models (LLMs) have emerged as a promising tool due to their ability to leverage rich semantic signals from textual side information (e.g., reviews, descriptions) alongside interaction logs. However, moving beyond simple pattern matching benefits from treating the LLM not merely as an autoregressive generator of item candidates or ranked lists from textualized user histories, but as an implicit probabilistic model over latent user intent, in line with emerging theoretical frameworks that interpret in-context learning as approximate Bayesian inference over latent variables (Bengio et al., 2025; Wang et al., 2023; Li et al., 2023). This motivates seeking explicit, time-varying preference distributions over interpretable semantic clusters that can be analyzed, audited, and reused across downstream applications.
The prevailing approach for capturing user preferences involves leveraging LLMs to directly generate ranked lists of items (Wang et al., 2025, 2024a; Ngo and Nguyen, 2024; Yu et al., 2024; Deng et al., 2025). Even recent methods adopting textual rationales or chain-of-thought ultimately rely on this autoregressive decoding process to produce final recommendations (Wang et al., 2024b; Yi et al., 2025). While effective for predicting short-term interactions with already-popular item, this paradigm leaves critical gaps. First, it keeps the latent preference representation implicit: generation fails to expose an explicit, structured posterior over preferences, making it infeasible for audits and explanations. Second, in direct generation, reliance on decoding heuristics (e.g., beam search, top-$k$ sampling) could exacerbate head concentration by truncating low-probability clusters. As detailed in Section˜3.3, this structural vulnerability to frequency bias and search truncation might suppresses the long tail (Holtzman et al., 2020; Stahlberg and Byrne, 2019), preventing faithful personalization of user-specific niche preferences.
These limitations have significant consequences for user-facing AI systems. The lack of an explicit user model hampers transparency and governance, making it difficult to systematically inspect or audit how preferences are represented across users and groups. At the same time, systematically suppressing low-probability but relevant items reinforces popularity during training and deployment, reducing content diversity and marginalizing long-tail preferences on real-world platforms (Bakshy et al., 2015; Wu et al., 2024). Together, these observations motivate a shift from sequence generation to explicit, distributional user modeling, leading to our central research question:
How can we infer personalized preference distributions via LLMs as explicit posteriors over semantic clusters that preserve diverse user preferences and support downstream decision-making?
To address this question, we propose SPECTRA (Semantic Preference Extraction and Clustered TRAcking), a distributional user modeling framework that treats the LLM as an implicit probabilistic model and extracts a cluster-level probability vector over a semantically interpretable set of preference clusters by probing the model’s logits. Using a cluster vocabulary (e.g., tags or categories) provides a stable semantic interface, avoiding the volatility and sparsity of item-level identifiers while yielding preference representations that are comparable across users and time. By representing preferences as distributions – rather than decoded item lists – SPECTRA reduces head-dominated concentration and better preserves niche preferences, improving coverage and exposure diversity in the resulting preference representation. For scalability to large cluster space, SPECTRA remains tractable via coarse-to-fine inference over a semantic taxonomy.
To validate SPECTRA, we evaluate on three datasets spanning varying scale and cluster complexity: MovieLens, Yelp, and a large-scale industrial short-video platform. Across these settings, SPECTRA consistently outperforms direct generation (DG) and strong LLM-based ranking baselines, improving distributional alignment, exposure diversity, and downstream ranking quality. Beyond short-horizon next-step prediction, our experiments further show that SPECTRA more reliably captures persistent, long-term preference structure over time, yielding stronger performance in long-horizon evaluation settings. SPECTRA achieves particularly significant gains on under-represented preference segments, consistent with its ability to mitigate head-dominated concentration; accordingly, SPECTRA is not intended to replace mature head-optimized production rankers, but to complement them by recovering user-specific niche preferences that are often suppressed under decoding-based pipelines. Our primary contributions are threefold:
@@
    Distributional user modeling from LLMs. We reframe LLM-based user modeling from autoregressive list generation to distributional inference, producing cluster-level preference distributions over a semantically interpretable cluster vocabulary.
@@
    Scalable posterior extraction via probing. We extract these cluster-level distributions by probing LLM logits, and design inference procedures that remain tractable for large cluster vocabularies via coarse-to-fine hierarchical inference over a semantic taxonomy.
@@
    Extensive Empirical Validation. Across MovieLens, Yelp, and a large-scale industrial short-video platform, SPECTRA yields consistent gains in distributional fidelity and coverage: it improves alignment to empirical preference distributions (e.g., $23\%$ JS divergence vs. a strong LLM-based ranking baseline such as Qwen3-Reranker-8B (QwenTeam, 2025)) and increases exposure diversity (up to $30\%$ global exposure entropy) relative to direct generation. As a downstream consequence, these representation gains also improve ranking outcomes (up to $40\%$ NDCG on public benchmarks) and deliver up to 7× higher NDCG on under-represented preference segments than a production Transformer baseline.
Figure 1: The SPECTRA framework pipeline. Left (Training): we adapt an LLM (CPT + SFT) to predict a user’s future preference distribution over $K$ clusters from historical context. Center (Inference): we extract cluster scores from the model’s logits via likelihood-based probing or generative classification (yielding score vector $S$, and obtain a predicted distribution $\hat{\theta}=\text{softmax}(S)$ (and ranking via $\text{argsort}(S)$). Application: the transparent distribution supports personalized ranking, long-tail interest discovery, and interpretable preference profiling.
@@ 2 Related Work
User Preference Representation. Prior work models user preferences from interaction histories using multi-preference representations (Shi et al., 2023; Cen et al., 2020; Zhou et al., 2018) and sequential next-item predictors (Rendle et al., 2010; Hidasi et al., 2015; Kang and McAuley, 2018; Sun et al., 2019). These methods typically encode user state implicitly (as embeddings or item-level scores), which limits transparency for analysis and auditing (Doshi-Velez and Kim, 2017; Lipton, 2018; Rudin, 2019). An open challenge is to surface an explicit, cluster-level preference distribution that is directly comparable across time.
LLMs for Preference Representation. LLMs have been incorporated into preference representation via prompt-based unification (P5; (Geng et al., 2022)) and text-aware representation learning (U-BERT; (Qiu et al., 2021)), enabling direct list generation and generative ranking (Ngo and Nguyen, 2024; Deng et al., 2025; Wang et al., 2024a, 2025) as well as textual profile generation for explainability (Yi et al., 2025). Recent work enriches this with review-grounded preference cues (Kim et al., 2025), user-embedding contextualization (Ning et al., 2025), and LLM-enhanced generative retrieval (Paischer et al., 2025). Reasoning-style pipelines (e.g., CoT4Rec, LLMRG) generate item-level rationales for recommendations (Yue et al., 2025; Wang et al., 2024b). These approaches typically keep user state implicit (in hidden representations or text) or operate at the item level. In contrast, SPECTRA addresses the upstream task of extracting a time-comparable, cluster-level preference distribution that can serve as a structured signal for downstream rankers or reasoning modules.
Belief Extraction and Scalable Inference. Belief extraction from token likelihoods has been studied mainly for factual knowledge probing and prompt-based classification (Petroni et al., 2019; Schick and Schütze, 2020; Raffel et al., 2020; Zhou et al., 2024). This paradigm can also be applied to infer preference distributions over semantic clusters in user modeling settings. Direct decoding can suffer from decoding/search artifacts (Holtzman et al., 2020; Stahlberg and Byrne, 2019), motivating inference procedures that avoid reliance on decoding heuristics. Distributional objectives in alignment and uncertainty modeling (Siththaranjan et al., 2023; Yao et al., 2024; Melnyk et al., 2024) further motivate predicting an explicit preference distribution. Scaling to large cluster spaces commonly uses coarse-to-fine hierarchical inference, as in extreme multi-label classification (Zhou et al., 2024).